OpenStudy (anonymous):

evaluate the integral lnx/(x*sqrt((lnx)^2+1) very confused

ganeshie8 (ganeshie8):

You cannot break a radical... So whenever u see radical, the first thing to try is to substitute the stuff inside radical

OpenStudy (anonymous):

yeah i set u = lnx^2+1

ganeshie8 (ganeshie8):

careful lnx^2 is very much different from (lnx)^2

OpenStudy (anonymous):

hmm this is what i got so far 1/2 integral 1/x * u^-1/2

OpenStudy (anonymous):

du/2=lnx

ganeshie8 (ganeshie8):

hold up

ganeshie8 (ganeshie8):

\(\large u = (\ln x)^2 + 1 \) \(\large du = 2 (\ln x) (\ln x)' dx\) \(\large du = 2 (\ln x) (\dfrac{1}{x}) dx \) \(\large \dfrac{du}{2} = \dfrac{\ln x}{x} dx \)

OpenStudy (anonymous):

oh chain rule right?

ganeshie8 (ganeshie8):

yes :)

OpenStudy (anonymous):

those always screw me over

ganeshie8 (ganeshie8):

plug them in the integrand now it simplifies nicely

ganeshie8 (ganeshie8):

lol they does every one in the start...

ganeshie8 (ganeshie8):

\[\large \int \dfrac{ \ln x}{x*\sqrt{(\ln x)^2+1}} dx\] after substitution becomes : \[\large \int \dfrac{ 1}{\sqrt{u}} \dfrac{du}{2}\]

ganeshie8 (ganeshie8):

which is trivial to integrate

OpenStudy (anonymous):

1/2 integral u^(-1/2) turns into 1/2 * 2u^(1/2) which then is u^1/2 then i plug u back in so it's sqrt((lnx)^2+1)?

ganeshie8 (ganeshie8):

looks good, and dont forget the arbitrary constant